It is estimated that t years from now the population of a particular city will be....
P = t^2 - 24t + 96,000
How many years from now will the population be 95,865?

Thanks in advance, I appreciate it!

Answer by jim_thompson5910(35256)

P = t^2 - 24t + 96000

95865 = t^2 - 24t + 96000

0 = t^2 - 24t + 96000 - 95865

0 = t^2 - 24t + 135


Now use the quadratic formula to solve for t:


t = (-b+-sqrt(b^2-4ac))/(2a)

t = (-(-24)+-sqrt((-24)^2-4(1)(135)))/(2(1))

t = (24+-sqrt(576-(540)))/(2)

t = (24+-sqrt(36))/2

t = (24+sqrt(36))/2 or t = (24-sqrt(36))/2

t = (24+6)/2 or t = (24-6)/2

t = 30/2 or t = 18/2

t = 15 or t = 9


So this means that in 15 or 9 years the population will be 95,865


Note: the population will reach 95,865 in 9 years first, dip below 95865 and then climb back up to 95,865 when t = 15

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim